OpenStudy (anonymous):

Evaluate the following expression using the given values: Find x - 3y if x = 3 and y = -2. -9 -3 3 9

jigglypuff314 (jigglypuff314):

plug in so if: x - 3y and x=3 and y=-2 then (3)-3*(-2) then you multiply and add

OpenStudy (anonymous):

x-3y is the given eqn...we need to put the values of x any y in their respective places.. 3-(3*(-2)) =3+6 =9

OpenStudy (anonymous):

Thank Youu. Soo Much I ahev ANother ONe I Got To Ask.

OpenStudy (anonymous):

have u undrstood dis one?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

ok ask